Find limit of sequence =n+2/3-n
1
Expert's answer
2020-05-03T17:31:55-0400

Let "(x_n)=\frac{n+2}{3-n}" be a sequence in "\R" .

"\therefore \lim_{n\to \infin} (x_n)=lim_{n\to \infin} \frac{n+2}{3-n}"

Divided numerator and denominator by n. We get


"lim_{n\to \infin}(x_n)=lim_{n\to \infin} \frac{1+\frac{2}{n}}{\frac{3}{n}-1}"

"=\frac{1}{-1}=-1"